OpenStudy (iwanttogotostanford):

Simplify open parentheses x to the 2 thirds power close parentheses to the 4 fifths power.

OpenStudy (iwanttogotostanford):

@Koikkara

OpenStudy (iwanttogotostanford):

@pooja195

OpenStudy (iwanttogotostanford):

choices: x to the 2 fifths power x to the 2 fifteenths power x to the 8 fifteenths power x to the 22 fifteenths power

zepdrix (zepdrix):

Exponent power rule:\[\large\rm (a^b)^c\quad= a^{b\cdot c}\]Simply multiply. \[\large\rm \left[x^{2/3}\right]^{4/5}\quad=x^{\frac{2}{3}\cdot\frac{4}{5}}\]